# Estimating the D-line Energy Splitting of Alkali Metals Using a Modified Perturbation Method

### v1

@@ 5 Conclusion
In this report we demonstrated how a well known result for the D line splitting energy in hydrogen can be modified and then used to estimate this values for all of the alkali metals. The alteration is in replacing the Bohr energy with the first ionization potential for the element involved. This outcome then lead us to find values for screening constants that could be used to describe this effect whilst using the hydrogenic model. The screening constant is found to be a mildly quadratic function of the atomic number. Further, they produce a smaller screening effect than the screening constants generated to predict ionization potentials, a discrepancy we attribute to the electron relaxation effect.
By relating the problem to a classical situation from electrostatics, we are able to write an approximate formula for the screening constant in terms of the ratio of cation and atomic radii. It is hoped that this result might find use in predicting fine structure splitting energies for other atomic groups from the periodic table.

### v2

@@ 5 Conclusion
In this report we demonstrated how a well known result for the D-line splitting energy in hydrogen can be modified and then used to estimate this value for all of the alkali metals. The alteration is in adjusting the Bohr energy with an appropriate screening constant which is a function of the atomic number. Through an empirical fit the screening constant is shown to be a mildly quadratic function of the atomic number. Using a simple model, we show how this form of the screening constant can be derived from first principles with a relativistic correction. It is hoped that this result might find use in predicting fine structure splitting energies for other atomic groups from the periodic table.